# Pontryagin Neural Operator for Solving General-Sum Differential Games with Parametric State Constraints

### v1

@@ 4 Pontryagin Neural Operator
Following standard treatment in neural operators, we introduce an input function $a(\cdot,\cdot):\mathcal{X}\times\Theta^{2}\rightarrow(0,1)$ to encode parametric settings of state constraints: $a(\textbf{x},\boldsymbol{\theta})=1$ if x violates the constraints according to $\boldsymbol{\theta}$, or otherwise $a(\textbf{x},\boldsymbol{\theta})=0$. Let $X\in\mathbb{R}^{N\times d_{x}}$ be a lattice of $\mathcal{X}$, we denote by $a(X,\boldsymbol{\theta})\in\{0,1\}^{N}$ the batch Boolean outputs at all $N$ lattice nodes. Then $\hat{\vartheta}$ is defined as a linear combination of function bases:
@@
where the branch network $\{b_{k}\}:\{0,1\}^{N}\rightarrow\mathbb{R}$ encodes the PDE parameters into function coefficients, and the trunk network $\{t_{k}\}:\mathbb{R}^{d_{x}}\times[0,T]\rightarrow\mathbb{R}$ encodes the input information into basis function values.
Figure 1: Illustration of Pontryagin Neural Operator. PNO extends DeepONet (Lu et al.(2021)Lu, Jin, Pang, Zhang, and Karniadakis) to Pontryagin-mode PINN. The value $\hat{\vartheta}$ is decomposed into basis functions (trunk) and their HJI-parameter-dependent coefficients (branch). We use losses on costate predictions to regularize the learning, which is key to the successful learning of values with large Lipschitz constants when state constraints are present. The baseline hybrid neural operator is similar to PNO, but without the learnable costate net. Instead, it solves BVPs as an overhead and use these fixed supervisory data for regularization.
@@
We introduce an additional costate network $\hat{\lambda}(\cdot,\cdot):\mathcal{X}\times[0,T]\rightarrow\mathcal{X}_{*}$, where $\mathcal{X}_{*}$ is the costate space. We can now evaluate the state, costate, and value trajectories starting from some sampled state and time $(\textbf{x},t)$ following their corresponding dynamics defined in Eq. (4), first by sequentially maximizing the Hamiltonian using $\hat{\lambda}$ to derive the state trajectory $\hat{\textbf{x}}\in\mathcal{X}^{K}$, and then using the transversality conditions of the costate and the value to derive the costate trajectory $\tilde{\lambda}\in\mathcal{X}_{*}^{K}$ and the value trajectory $\tilde{\vartheta}\in\mathbb{R}^{K}$ backward in time, where $K=\frac{T-t}{\Delta t}$ and $\Delta t$ is a small time interval. Given a batch of $N$ initial states, we collect forward state trajectories $\hat{\mathcal{X}}:=\{\hat{\textbf{x}}^{(n)}\}_{n=1}^{N}$, forward and backward costate trajectories, $\hat{\Lambda}:=\{\hat{\lambda}^{(n)}\}_{n=1}^{N}$ and $\tilde{\Lambda}:=\{\tilde{\lambda}^{(n)}\}_{n=1}^{N}$ respectively, and backward value trajectories $\tilde{\mathcal{V}}:=\{\tilde{\vartheta}^{(n)}\}_{n=1}^{N}$. Lastly, we denote the conventional PINN dataset for formulating the PDE and boundary losses as $\mathcal{D}:=\{(\textbf{x},t)^{(n)}\in\mathcal{X}\times[0,T]\}_{n=1}^{N_{L}}$ and $\mathcal{D}_{D}:=\{\textbf{x}^{(n)}\in\mathcal{X}\}_{n=1}^{N_{D}}$. The PNO training loss with respect to value and costate functions of both players $(\hat{\boldsymbol{\vartheta}},\hat{\boldsymbol{\lambda}})$ is then defined as:
$L_{PNO}(\hat{\boldsymbol{\vartheta}},\hat{\boldsymbol{\lambda}}):=$ $\sum_{i=1}^{2}\sum_{\mathcal{D}}\left\|L(\hat{\boldsymbol{\vartheta}}_{i}^{(n)},\nabla_{\textbf{x}_{i}}\hat{\vartheta}_{i}^{(n)},\textbf{x}_{i}^{(n)},t^{(n)})\right\|_{1}+\sum_{\mathcal{D}_{D}}C_{1}\left\|D(\hat{\vartheta}_{i}^{(n)},\textbf{x}_{i}^{(n)})\right\|_{1}$ (6)
@@
where $\hat{\vartheta}_{i}^{(n)}$, $\tilde{\vartheta}_{i}^{(n)}$, $\hat{\lambda}_{i}^{(n)}$, $\tilde{\lambda}_{i}^{(n)}$ are abbreviations for $\hat{\vartheta}(\textbf{x}_{i}^{(n)},t^{(n)},\boldsymbol{\theta}_{i}^{(n)})$, $\tilde{\vartheta}_{i}(\textbf{x}_{i}^{(n)},t^{(n)},\boldsymbol{\theta}_{i}^{(n)})$, $\hat{\lambda}_{i}(\textbf{x}_{i}^{(n)},t^{(n)})$, $\tilde{\lambda}(\textbf{x}_{i}^{(n)},t^{(n)})$ respectively. $C_{i}>0$ for $i=1,..,5$ are hyperparameters that balance the loss terms. We summarize the training of PNO in Alg. 5. Note that for the Pontryagin-mode losses to be differentiable with respect to $(\hat{\boldsymbol{\vartheta}},\hat{\boldsymbol{\lambda}})$, we use costate net and ODE solver (RK45) to compute the forward and backward trajectories $(\hat{\mathcal{X}},\hat{\Lambda},\tilde{\Lambda},\tilde{\mathcal{V}})$. Line 9-12 in Alg. 5 outline the computation process.
@@
For standard PINN residual and boundary datasets $\mathcal{D}$ and $\mathcal{D}_{D}$, we use a curriculum learning scheme following (Bansal and Tomlin(2021); Krishnapriyan et al.(2021)Krishnapriyan, Gholami, Zhe, Kirby, and Mahoney) to observe temporal causality of value functions. For Pontryagin-mode trajectories $(\hat{\mathcal{X}},\hat{\Lambda},\tilde{\Lambda},\tilde{\mathcal{V}})$, our experimental results show that an importance sampling strategy is critical. We start by sampling $N$ initial states uniformly in $\mathcal{X}$. Subsequently, we follow the evolutionary sampling algorithm (Daw et al.(2022)Daw, Bu, Wang, Perdikaris, and Karpatne) to compute absolute values of the PDE residual $r(\textbf{x}_{i}^{(n)})=\|L(\boldsymbol{\hat{\vartheta}}_{i}^{(n)},\nabla_{\textbf{x}_{i}^{(n)}}\hat{\vartheta}_{i}^{(n)},\textbf{x}_{i}^{(n)},t^{(n)})\|_{1}$ for $n=1,...,N$ using Eq. (3). We iteratively sample and remove samples with residuals lower than the average. The resultant batch represents initial states for which the current $\hat{\boldsymbol{\vartheta}}$ cannot generalize well. The sampled initial states gradually accumulate in the region with informative collision and near-collision knowledge during the training. Convergence of the evolutionary-based PINN has been proved in (Daw et al.(2022)Daw, Bu, Wang, Perdikaris, and Karpatne).
@@
PNO does not require offline data collection through solving BVPs and thus avoids typically convergence issues encountered in solving differential games with multiple local equilibria and slow convergence dynamics. While a full PAC learning proof of PNO is not yet available, our results suggest that PNO achieves more data-efficient learning of values and policies than an existing hybrid method that takes advantage of informative samples that reveal structures of the value or costate landscapes. One hypothesis for the success of PNO is that it decomposes the value landscape into (terminal) boundaries and value dynamics along trajectories (a Lagrangian perspective), i.e., the hard-to-learn value with large Lipschitz constant can potentially be learned more effectively when each of its components, namely, the continuous boundary, continuous trajectories, and discontinuous value dynamics, are easy to learn given the others. We note that the improvement in data-efficiency of PINN through the method of characteristics has been discussed in (Mojgani et al.(2023)Mojgani, Balajewicz, and Hassanzadeh) for 2D convection-diffusion and Burger’s equations, but not yet in the context of high-dimensional HJ equations for optimal control or differential games.
@@ 5 Experiments and Results
To demonstrate the efficacy of PNO, we solve HJI equations for the interaction between two vehicles at an uncontrolled intersection. Each vehicle has two state variables (location and velocity) and one parameter that defines their perception of the collision zone. Hence we have a pair of 5D HJI equations defined in a 2D parameter space. We compare PNO with the hybrid method from (Zhang et al.(2023a)Zhang, Ghimire, Zhang, Xu, and Ren) using both sample equilibrial trajectories solved from BVP (Eq. (4)) and the value landscape solved by a general-sum extension of the level set algorithm (Bui et al.(2022)Bui, Giovanis, Chen, and Shriraman) (which cannot be scaled to larger problems). Due to space limitation, we will focus on closed-loop safety performance as the performance metric. Hardware. Supervised equilibria data and closed-loop intersection simulations are computed on a workstation with 3.50GHz Xeon E5-1620 v4 and one GTX TITAN X with 12 GB memory. Hybrid and Pontryagin neural operators are trained on an A100 GPU with 40 GB memory.
@@
$\hat{\Lambda}\leftarrow$ compute forward costate trajectories using $\hat{\lambda}$
$\tilde{\Lambda}\leftarrow$ compute backward costate trajectories as solutions to inverse value problems with terminal values $\hat{\mathcal{X}}(T)$ and $\hat{\Lambda}(T)$ via solving Eq. (4) using ODE solver (RK45)
$\tilde{\mathcal{V}}\leftarrow$ compute value trajectories with $\hat{\mathcal{X}}$ and $\tilde{\Lambda}$ via solving Eq. (1) and Eq. (4)
@@
We adopt settings of the uncontrolled intersection case studied in (Zhang et al.(2023a)Zhang, Ghimire, Zhang, Xu, and Ren): Each vehicle is represented by states $x_{i}:=(d_{i},v_{i})$, $d_{i}$ for location and $v_{i}$ for velocity. Both vehicles follow simple dynamics: $\dot{d}_{i}=v_{i}$ and $\dot{v}_{i}=u_{i}$, where $u_{i}\in[-5,10]m/s^{2}$ is the control input. Let $R$, $L$, and $W$ be the road length, vehicle length, and vehicle width, respectively. The instantaneous loss and state constraint are defined as:
@@
$\sigma(d,\theta)=$ $\left(1+\exp(-\gamma(d-R/2+\theta W/2))\right)^{-1}$ (8)
$\left(1+\exp(\gamma(d-(R+W)/2-L))\right)^{-1}.$
@@
Player Types Learning Method Metrics Ground Truth  | Hybrid Neural Operator  | Pontryagin Neural Operator (1, 1) (1, 2) (1, 3) (1, 4) (1, 5) 0.00% 3.00% 0.33% 2.67% 21.8% 6.50% 4.83% 9.50% 9.17% 9.33% 9.83% 8.00% 9.00% 8.00% 8.17% (2, 1) (2, 2) (2, 3) (2, 4) (2, 5) 2.67% 21.8% 6.50% 7.00% 34.7% 8.33% 14.0% 26.3% 15.2% 21.5% 36.2% 21.5% 28.7% 41.5% 26.8% ($\theta_{1}$, $\theta_{2}$) (3, 1) (3, 2) (3, 3) (3, 4) (3, 5) Col.$\downarrow$ 4.83% 9.50% 9.17% 14.0% 26.3% 15.2% 23.7% 22.5% 23.5% 33.3% 32.7% 33.2% 43.0% 41.5% 41.3% (4, 1) (4, 2) (4, 3) (4, 4) (4, 5) 9.33% 9.83% 8.00% 21.5% 36.2% 21.5% 33.3% 32.7% 33.2% 42.3% 40.2% 41.3% 54.3% 51.8% 52.3% (5, 1) (5, 2) (5, 3) (5, 4) (5, 5) 9.00% 8.00% 8.17% 28.7% 41.5% 26.8% 43.0% 41.5% 41.3%54.3% 51.8% 52.3% 61.0% 60.0% 60.2%
@@
To demonstrate the advantage of PNO, we perform a comparison in favor of the baseline (the hybrid neural operator): For the baseline, we generate 1k ground truth trajectories from initial states uniformly sampled in $\mathcal{X}_{GT}:=[15,20]m\times[18,25]m/s$ by solving Eq. (4). This set is identified to contain a significant amount of corner cases (e.g., $61\%$ of $\mathcal{X}_{GT}$ yield collisions when $\boldsymbol{\theta}=(5,5)$, see Tab. 1) and will also be used for test. Each trajectory consists of 31 data points with a time interval of $0.1s$ for each of the players, resulting in a total of 62k data points. For PNO, the costate net explores the initial state within $\mathcal{X}_{HJ}:=[15,105]m\times[15,32]m/s$ and generates 1k closed-loop trajectories including 62k data points. $\mathcal{X}_{HJ}$ is much less informed than $\mathcal{X}_{GT}$. To improve training efficiency, we eliminate data points generated by the costate net that fall beyond the state bounds and normalize the input data to $[-1,1]$ for the trunk and costate net of both neural operators. Additionally, the trunk net uniformly samples 60k states in $\mathcal{X}_{HJ}:=[15,105]m\times[15,32]m/s$ for model refinement. Lastly, we train the model using four player-type configurations $(\theta_{1},\theta_{2})=\{(1,1),(1,5),(5,1),(5,5)\}$ and evaluate the model generalization performance within the type space $\Theta$. We reiterate that the total number of data points for the two methods are approximately the same.
@@
Player Types Learning Method Metrics Ground Truth  | Hybrid Neural Operator  | Pontryagin Neural Operator (1, 1) (1, 2) (1, 3) (1, 4) (1, 5) 0.00% 3.00% 0.33% 0.00% 19.5% 4.33% 0.00% 6.33% 5.67% 0.00% 2.17% 0.00% 0.00% 0.83% 1.00% (2, 1) (2, 2) (2, 3) (2, 4) (2, 5) 0.00% 19.5% 4.33% 0.00% 30.0% 2.83% 0.00% 14.7% 3.83% 0.00% 19.0% 1.17% 0.00% 20.2% 0.00% ($\theta_{1}$, $\theta_{2}$) (3, 1) (3, 2) (3, 3) (3, 4) (3, 5) Col.$\downarrow$ 0.00% 6.33% 5.67% 0.00% 14.7% 3.83% 0.00% 0.50% 0.67% 0.00% 1.00% 0.50% 0.00% 1.17% 0.67% (4, 1) (4, 2) (4, 3) (4, 4) (4, 5) 0.00% 2.17% 0.00% 0.00% 19.0% 1.17% 0.00% 1.00% 0.50% 0.00% 0.00% 0.00% 0.00% 0.33% 0.33% (5, 1) (5, 2) (5, 3) (5, 4) (5, 5) 0.00% 0.83% 1.00% 0.00% 20.2% 0.00% 0.00% 1.17% 0.67% 0.00% 0.33% 0.33% 0.00% 0.00% 0.00%
Figure 2: (a, d): Ground truth trajectories (projected to $d_{1}$-$d_{2}$) w and w/o inevitable collisions, respectively. (b-c), (e-f): Trajectories generated using hybrid and PNO w and w/o inevitable collisions, respectively. Color: Actual values of Player 1. Red dots: Initial states that yield collisions. Solid gray box: collision area for player type $\boldsymbol{\theta}=(1,1)$.
@@
The neural operators utilize fully-connected networks with 3 hidden layers, each comprising 64 neurons with tanh activation. The hybrid neural operator uses the Adam optimizer with a fixed learning rate of $2\times 10^{-5}$ for pre-training the network over 100k iterations using supervised data. It then combines the supervised data with states sampled from an expanding time window, starting from the terminal time, to refine the model for an additional 100k iterations. PNO uses the Adam optimizer with an adaptive learning rate, starting from $2\times 10^{-5}$, and is initially trained to satisfy the boundary conditions over 50k iterations. Subsequently, the network undergoes refinement through 3k gradient steps per epoch, encompassing a total of 10 epochs for every 10 training iterations. The costate net resamples the initial state to generate closed-loop trajectories every 10 epochs, and these data points are used to train the model. The overall network refinement process spans 300 training iterations. Both PNO and hybrid incorporate adaptive activation functions (Jagtap et al.(2020)Jagtap, Kawaguchi, and Karniadakis). In terms of wall-clock computational cost, the hybrid neural operator requires approximately 37 hours, including 5.5 hours for data acquisition and 31.5 hours for model training. PNO requires 27 hours for model training. For comparison, solving all 25 individual HJI PDEs in $\Theta^{2}$ with an acceptable meshgrid resolution using an efficient level set package (Bui et al.(2022)Bui, Giovanis, Chen, and Shriraman) would require $>50$ hours. The resultant value approximation does not guarantee safety performance since value gradient approximation error is not controlled.

### v2

@@ 4 Pontryagin Neural Operator
Following standard treatment in neural operators, we introduce an input function $a(\cdot,\cdot):\mathcal{X}\times\Theta^{2}\rightarrow(0,1)$ to encode parametric settings of state constraints: $a(\textbf{x},\boldsymbol{\theta})=1$ if x violates the constraints according to $\boldsymbol{\theta}$, or otherwise $a(\textbf{x},\boldsymbol{\theta})=0$. Let $X\in\mathbb{R}^{L\times d_{x}}$ be a lattice of $\mathcal{X}$, we denote by $a(X,\boldsymbol{\theta})\in\{0,1\}^{L}$ the batch Boolean outputs at all $L$ lattice nodes. Then $\hat{\vartheta}$ is defined as a linear combination of function bases:
@@
where the branch network $\{b_{k}\}:\{0,1\}^{L}\rightarrow\mathbb{R}$ encodes the PDE parameters into function coefficients, and the trunk network $\{t_{k}\}:\mathbb{R}^{d_{x}}\times[0,T]\rightarrow\mathbb{R}$ encodes the input information into basis function values.
Figure 1: Illustration of Pontryagin Neural Operator. PNO extends DeepONet (Lu et al., 2021) to Pontryagin-mode PINN. The value $\hat{\vartheta}$ is decomposed into basis functions (trunk) and their HJI-parameter-dependent coefficients (branch). We use losses on costate predictions to regularize the learning, which is key to the successful learning of values with large Lipschitz constants when state constraints are present. The baseline hybrid neural operator is similar to PNO, but without the learnable costate net. Instead, it solves BVPs as an overhead and use these fixed supervisory data for regularization.
@@
We introduce an additional costate network $\hat{\lambda}(\cdot,\cdot):\mathcal{X}\times[0,T]\rightarrow\mathcal{X}_{*}$, where $\mathcal{X}_{*}$ is the costate space. We can now evaluate the state, approximate costate, and approximate value trajectories starting from some sampled state and time $(\textbf{x},t)$ following their corresponding dynamics defined in Eq. (4), first by sequentially maximizing the Hamiltonian using forward approximate $\hat{\lambda}$ to derive the state trajectory $\hat{\textbf{x}}\in\mathcal{X}^{K}$, and then using the transversality conditions of the approximate costate and value to derive approximate costate trajectory $\tilde{\lambda}\in\mathcal{X}_{*}^{K}$ and approximate value trajectory $\tilde{\vartheta}\in\mathbb{R}^{K}$ backward in time, where $K=\frac{T-t}{\Delta t}$ and $\Delta t$ is a small time interval. Given a batch of $N$ initial states, we collect forward approximate state trajectories $\hat{\mathcal{X}}:=\{\hat{\textbf{x}}^{(n)}\}_{n=1}^{N}$, forward and backward approximate costate trajectories, $\hat{\Lambda}:=\{\hat{\lambda}^{(n)}\}_{n=1}^{N}$ and $\tilde{\Lambda}:=\{\tilde{\lambda}^{(n)}\}_{n=1}^{N}$ respectively, and backward approximate value trajectories $\tilde{\mathcal{V}}:=\{\tilde{\vartheta}^{(n)}\}_{n=1}^{N}$. Lastly, we denote the conventional PINN dataset for formulating the PDE and boundary losses as $\mathcal{D}:=\{(\textbf{x},t)^{(n)}\in\mathcal{X}\times[0,T]\}_{n=1}^{N_{L}}$ and $\mathcal{D}_{D}:=\{\textbf{x}^{(n)}\in\mathcal{X}\}_{n=1}^{N_{D}}$. The PNO training loss with respect to value and costate functions of both players $(\hat{\boldsymbol{\vartheta}},\hat{\boldsymbol{\lambda}})$ is then defined as:
$L_{PNO}(\hat{\boldsymbol{\vartheta}},\hat{\boldsymbol{\lambda}}):=$ $\sum_{i=1}^{2}\sum_{\mathcal{D}}\left\|L(\hat{\vartheta}_{i}^{(n)},\nabla_{\textbf{x}_{i}}\hat{\vartheta}_{i}^{(n)},\textbf{x}_{i}^{(n)},t^{(n)})\right\|_{1}+\sum_{\mathcal{D}_{D}}C_{1}\left\|D(\hat{\vartheta}_{i}^{(n)},\textbf{x}_{i}^{(n)})\right\|_{1}$ (6)
@@
where $\hat{\vartheta}_{i}^{(n)}$, $\tilde{\vartheta}_{i}^{(n)}$, $\hat{\lambda}_{i}^{(n)}$, $\tilde{\lambda}_{i}^{(n)}$ are abbreviations for $\hat{\vartheta}(\textbf{x}_{i}^{(n)},t^{(n)},\boldsymbol{\theta}_{i}^{(n)})$, $\tilde{\vartheta}_{i}(\textbf{x}_{i}^{(n)},t^{(n)},\boldsymbol{\theta}_{i}^{(n)})$, $\hat{\lambda}_{i}(\textbf{x}_{i}^{(n)},t^{(n)})$, $\tilde{\lambda}(\textbf{x}_{i}^{(n)},t^{(n)})$ respectively. $C_{i}>0$ for $i=1,..,5$ are hyperparameters that balance the loss terms. We summarize the training of PNO in Alg. 5. Note that for the Pontryagin-mode losses to be differentiable with respect to $(\hat{\boldsymbol{\vartheta}},\hat{\boldsymbol{\lambda}})$, we use costate net and ODE solver (RK45) to compute the forward and backward trajectories $(\hat{\mathcal{X}},\hat{\Lambda},\tilde{\Lambda},\tilde{\mathcal{V}})$. Line 9-12 in Alg. 5 outline the computation.
@@
For standard PINN residual and boundary datasets $\mathcal{D}$ and $\mathcal{D}_{D}$, we use a curriculum learning scheme following Bansal and Tomlin (2021); Krishnapriyan et al. (2021), where states are sampled from an expanding time window starting from the terminal time, to observe temporal causality of value functions. For Pontryagin-mode trajectories $(\hat{\mathcal{X}},\hat{\Lambda},\tilde{\Lambda},\tilde{\mathcal{V}})$, our experimental results show that an importance sampling strategy is critical. We start by sampling $N$ initial states uniformly in $\mathcal{X}$. Subsequently, we follow the evolutionary sampling algorithm (Daw et al., 2022) to compute absolute values of the PDE residual $r(\textbf{x}_{i}^{(n)})=\|L(\hat{\vartheta}_{i}^{(n)},\nabla_{\textbf{x}_{i}^{(n)}}\hat{\vartheta}_{i}^{(n)},\textbf{x}_{i}^{(n)},t^{(n)})\|_{1}$ for $n=1,...,N$ using Eq. (3) [^1] Our empirical studies showed that evolution based on costate losses achieves similar safety performance. Analytical understanding about the sampling-performance relation is yet to be established.. We iteratively sample and remove samples with residuals lower than the average. The resultant batch represents initial states for which the current $\hat{\boldsymbol{\vartheta}}$ cannot generalize well. The sampled initial states gradually accumulate in the region with informative collision and near-collision knowledge during the training. Convergence of the evolutionary-based PINN has been proved in Daw et al. (2022).
@@
PNO does not require offline data collection through solving BVPs and thus avoids typical convergence issues encountered in solving differential games with multiple local equilibria and slow convergence dynamics. While a full probably approximately correct (PAC) learning proof of PNO is not yet available, our results suggest that PNO achieves more data-efficient learning of values and policies than an existing hybrid method that takes advantage of informative samples that reveal structures of the value or costate landscapes. One hypothesis for its success is that PNO decomposes the value landscape into (terminal) boundaries and value dynamics along trajectories (a Lagrangian perspective), i.e., the hard-to-learn value with large Lipschitz constant can potentially be learned more effectively when each of its components, namely, the continuous boundary, continuous trajectories, and discontinuous value dynamics, are easy to learn given the others. We note that the improvement in data-efficiency of PINN through the method of characteristics has been discussed in Mojgani et al. (2023) for 2D convection-diffusion and Burger’s equations, but not yet in the context of high-dimensional HJ equations for optimal control or differential games.
@@ 5 Experiments and Results
To demonstrate the efficacy of PNO, we solve HJI equations for the interaction between two vehicles at an uncontrolled intersection. Each vehicle has two state variables (location $d_{i}$ and velocity $v_{i}$) and one parameter (player type $\theta$) that defines their perception of the collision zone. Hence we have a pair of 5D HJI equations defined in a 2D parameter space. We compare PNO with the hybrid method from Zhang et al. (2023a) using both sample equilibrial trajectories solved from BVP (Eq. (4)) and the value landscape solved by a general-sum extension of the level set algorithm (Bui et al., 2022), which cannot be scaled to larger problems. Due to space limitation, we will focus on closed-loop safety performance as the performance metric. Hardware. Supervised equilibria data and closed-loop intersection simulations are computed on a workstation with 3.50GHz Xeon E5-1620 v4 and one GTX TITAN X with 12 GB memory. Hybrid and Pontryagin neural operators are trained on an A100 GPU with 40 GB memory.
@@
$\hat{\Lambda}\leftarrow$ compute forward approximate costate trajectories using $\hat{\lambda}$
$\tilde{\Lambda}\leftarrow$ compute backward approximate costate trajectories as solutions to inverse value problems with terminal values $\hat{\mathcal{X}}(T)$ and $\hat{\Lambda}(T)$ via solving Eq. (4) using an ODE solver (RK45)
$\tilde{\mathcal{V}}\leftarrow$ compute approximate value trajectories with $\hat{\mathcal{X}}$ and $\tilde{\Lambda}$ via solving Eq. (1) and Eq. (4)
@@
We adopt settings of the uncontrolled intersection case studied in Zhang et al. (2023a): Each vehicle is represented by states $x_{i}:=(d_{i},v_{i})$. Both vehicles follow dynamics: $\dot{d}_{i}=v_{i}$ and $\dot{v}_{i}=u_{i}$, where $u_{i}\in[-5,10]m/s^{2}$ is the control input. Let $R$, $L$, and $W$ be the road length, vehicle length, and vehicle width, respectively. The instantaneous loss and state constraint are defined as:
@@
$\sigma(d,\theta)=\left(1+\exp(-\gamma(d-R/2+\theta W/2))\right)^{-1}\left(1+\exp(\gamma(d-(R+W)/2-L))\right)^{-1}.$ (8)
@@
Player Types Learning Method Metrics Ground Truth  | Hybrid Neural Operator  | Pontryagin Neural Operator (1, 1) (1, 2) (1, 3) (1, 4) (1, 5) 0.00% 0.17% 0.17% 2.67% 19.5% 6.50% 4.83% 10.2% 9.17% 9.33% 8.83% 8.00% 9.00% 7.50% 8.17% (2, 1) (2, 2) (2, 3) (2, 4) (2, 5) 2.67% 19.5% 6.50% 7.00% 24.2% 8.33% 14.0% 23.8% 15.2% 21.5% 26.5% 21.5% 28.7% 30.2% 26.8% ($\theta_{1}$, $\theta_{2}$) (3, 1) (3, 2) (3, 3) (3, 4) (3, 5) Col.$\downarrow$ 4.83% 10.2% 9.17% 14.0% 23.8% 15.2% 23.7% 26.5% 23.5% 33.3% 32.3% 33.2% 43.0% 41.0% 41.3% (4, 1) (4, 2) (4, 3) (4, 4) (4, 5) 9.33% 8.83% 8.00% 21.5% 26.5% 21.5% 33.3% 32.3% 33.2% 42.3% 40.3% 41.3% 54.3% 51.6% 52.3% (5, 1) (5, 2) (5, 3) (5, 4) (5, 5) 9.00% 7.50% 8.17% 28.7% 30.2% 26.8% 43.0% 41.0% 41.3% 54.3% 51.6% 52.3% 61.0% 59.7% 60.2%
@@
To demonstrate the advantage of PNO, we perform a comparison in favor of the baseline (the hybrid neural operator): For the baseline, we generate 1k ground truth trajectories from initial states uniformly sampled in an informed state set $\mathcal{X}_{GT}:=[15,20]m\times[18,25]m/s$ by solving Eq. (4). This set is identified to contain a significant amount of corner cases (e.g., $61\%$ of $\mathcal{X}_{GT}$ yield collisions when $\boldsymbol{\theta}=(5,5)$, see Tab. 1) and will also be used for test. Each trajectory consists of 31 data points with a time interval of $0.1s$ for each of the players, resulting in a total of 62k data points. For PNO, the costate net explores the initial state within a larger state set $\mathcal{X}_{HJ}:=[15,105]m\times[15,32]m/s$ and generates 1k closed-loop trajectories including 62k data points. $\mathcal{X}_{HJ}$ is much less informed than $\mathcal{X}_{GT}$ (less collisions). To improve training efficiency, we eliminate data points generated by the costate net that fall beyond the state bounds and normalize the input data to $[-1,1]$ for the trunk and costate net of both neural operators. Additionally, the trunk net uniformly samples 60k states in $\mathcal{X}_{HJ}:=[15,105]m\times[15,32]m/s$ for model refinement. Lastly, we train the model using four player-type configurations $(\theta_{1},\theta_{2})=\{(1,1),(1,5),(5,1),(5,5)\}$ and evaluate the model generalization performance within the type space $\Theta$. We reiterate that the total number of data points for the two methods are approximately the same.
@@
Player Types Learning Method Metrics Ground Truth  | Hybrid Neural Operator  | Pontryagin Neural Operator (1, 1) (1, 2) (1, 3) (1, 4) (1, 5) 0.00% 0.17% 0.17% 0.00% 17.0% 4.33% 0.00% 6.33% 5.67% 0.00% 1.00% 0.00% 0.00% 0.50% 1.00% (2, 1) (2, 2) (2, 3) (2, 4) (2, 5) 0.00% 17.0% 4.33% 0.00% 18.8% 2.83% 0.00% 11.3% 3.83% 0.00% 7.17% 1.17% 0.00% 4.50% 0.00% ($\theta_{1}$, $\theta_{2}$) (3, 1) (3, 2) (3, 3) (3, 4) (3, 5) Col.$\downarrow$ 0.00% 6.33% 5.67% 0.00% 11.3% 3.83% 0.00% 5.33% 0.67% 0.00% 0.33% 0.50% 0.00% 0.83% 0.67% (4, 1) (4, 2) (4, 3) (4, 4) (4, 5) 0.00% 1.00% 0.00% 0.00% 7.17% 1.17% 0.00% 0.33% 0.50% 0.00% 0.00% 0.00% 0.00% 0.33% 0.33% (5, 1) (5, 2) (5, 3) (5, 4) (5, 5) 0.00% 0.50% 1.00% 0.00% 4.50% 0.00% 0.00% 0.83% 0.67% 0.00% 0.33% 0.33% 0.00% 0.00% 0.00%
Figure 2: (a, d): Ground truth trajectories (projected to $d_{1}$-$d_{2}$) w and w/o inevitable collisions, respectively. (b-c), (e-f): Trajectories generated using hybrid and PNO w and w/o inevitable collisions, respectively. Color: Actual values of Player 1. Red dots: Initial states that yield collisions. Solid gray box: collision area for player type $\boldsymbol{\theta}=(1,1)$. Players with larger type parameters (e.g., $\boldsymbol{\theta}=(2,2)$) have correspondingly larger collision zones, as indicated by white box.
@@
The neural operators utilize fully-connected networks with 3 hidden layers, each comprising 64 neurons with tanh activation. The hybrid neural operator uses the Adam optimizer with a fixed learning rate of $2\times 10^{-5}$ for pre-training the network over 100k iterations using supervised data. It then combines the supervised data with states sampled from an expanding time window, starting from the terminal time, to refine the model for an additional 200k iterations. PNO uses the Adam optimizer with an adaptive learning rate, starting from $2\times 10^{-5}$, and is initially trained to satisfy the boundary conditions over 50k iterations. Subsequently, the network undergoes refinement through 3k gradient steps per epoch, encompassing a total of 10 epochs for every 10 training iterations. The costate net resamples the initial state to generate closed-loop trajectories every 10 epochs, and these data points are used to train the model. The overall network refinement process spans 300 training iterations. Both PNO and hybrid incorporate adaptive activation functions (Jagtap et al., 2020). In terms of wall-clock computational cost, the hybrid neural operator requires approximately 37 hours, including 5.5 hours for data acquisition and 31.5 hours for model training. PNO requires 27 hours for model training. For comparison, solving all 25 individual HJI PDEs in $\Theta^{2}$ with an acceptable meshgrid resolution using an efficient level set package (Bui et al., 2022) would require $>50$ hours. The resultant value approximation does not guarantee safety performance since value gradient approximation error is not controlled.